OpenStudy (anonymous):

In △JKL, solve for x. 17.32 38.16 66.73 74.89

OpenStudy (anonymous):

where is △JKL?

OpenStudy (anonymous):

using SOHCAHTOA the triangle has an opposite and an adjacent, so we use tan tan 27degree = 34/x

OpenStudy (anonymous):

Since you are given two sides corresponding to the angle 27 (opposite and adjacent), you can use the trigonometric ratio for tan of that angle to solve for x. tan(27) = 34/x => x = 34/tan(27). Use a calculator to find the value of x.

OpenStudy (anonymous):

so i would put tan(34?

OpenStudy (anonymous):

hmmm....I showed you that x = 34/tan(27). So, you should find tan(27) first and then divide 34 by that value to find x.

OpenStudy (anonymous):

got it thanks